OpenStudy (anonymous):

Last question I have before my final: How do I find tan(5pi/6) I would prefer u show the steps.

Nnesha (nnesha):

calculator ?? hmm or no ?

OpenStudy (anonymous):

Not that you really need one though.

Nnesha (nnesha):

well for final exam you really need one TI-84 calculator :D jk

OpenStudy (loser66):

\(tan (5\pi/6)=\dfrac{sin(5\pi/6)}{cos(5\pi/6)}\) And we know that \(sin(5\pi/6) = sin (\pi/6)=1/2\) \(cos (5\pi/6) = -cos(\pi/6)=-\dfrac{\sqrt3}{2}\) put them in and find tan (5pi/6)

Nnesha (nnesha):

yep but still you have to memorize unit circle or use 30-60-90 triangle theorem gO_Od luck!